### v2

# Diagrammatic Monte Carlo for finite systems at zero temperature

Abstract
High-order virtual excitations play an important role in microscopic models of nuclear reactions at intermediate energies. However, the factorial growth of their complexity has prevented their consistent inclusion in ab initio many-body calculations. For infinite systems at finite temperature, such drawbacks can be overcome using diagrammatic Monte Carlo (DiagMC) techniques to resum entire series of Feynman diagrams. We present a DiagMC algorithm that can be applied to self-bound systems with discrete energy levels at zero temperature, and demonstrate its potential for the Richardson model of nuclear pairing. We show that sampling the topological space of diagrams allows the inclusion of high-order excitations that are neglected in state-of-the-art approximations used in nuclear physics and quantum chemistry. We propose that sampling the diagrammatic space can overcome the long-standing gap between our microscopic understanding of structure and reactions in nuclear physics.
Ab initio nuclear theory has significantly advanced over the past two decades by combining effective field theories (EFT) of the nuclear force with the latest computational approaches to the quantum many-body problem. Several high-precision methods can predict the ground state properties of medium-light nuclei [1, 2], and recently reached heavy elements [3, 4]. The EFT interactions used in modern literature are rooted in the symmetries of QCD and avoid relying on the phenomenology of stable nuclei, so that controlled and accurate predictions remain feasible even toward the limits of nuclear existence. Exotic neutron-rich nuclei far from stability are essential for studying many fundamental open problems of modern physics, such as the evolution of the shell structure and the formation of heavy elements [5, 6]. In particular, the role of unstable nuclei in the astrophysical r-process in neutron star mergers cannot be overestimated [7, 8]. State-of-the-art and future facilities that work with unstable beams rely on nuclear reaction models to discover new isotopes and interpret experimental results [9, 10, 11, 12, 13, 14, 15]. Despite the success of the structure calculations discussed above, the microscopic description of nuclear reactions is more challenging [16]. Among the available ab initio methods [17, 18, 19, 20, 21], the self-consistent Green’s function (SCGF) approach is suited for a unified description of structure and reactions because of the wealth of physical information contained. The exact spectral distribution (which gathers all one-nucleon addition and removal spectroscopy), the ground-state energy and density profiles, and the exact microscopic optical potential are consistently included in the formalism  [22, 23, 24, 25]. The Green’s function $G$ is the exact solution of the Dyson equation $G_{\alpha\beta}(\omega)=G^{(0)}_{\alpha\beta}(\omega)+\sum_{\gamma\delta}G^{(0)}_{\alpha\gamma}(\omega)\Sigma^{\star}_{\gamma\delta}(\omega)G_{\delta\beta}(\omega)$, where $G^{(0)}$ describes a freely propagating particle (or hole) and the term $G^{(0)}\Sigma^{\star}G$ introduces recursive dynamic interactions with the other particles of the system. In this picture, the self-energy $\Sigma^{\star}$ is a complex, nonlocal, energy-dependent microscopic optical potential [26, 22, 23]. SCGF theory relies on systematically improvable many-body truncations of the self-energy [27], however, current approximation schemes do not include high-order configurations that dominate mid-energy scattering [19]. Due to the factorial growth of the Feynman diagrams (see Fig. 1), improving the accuracy of ab initio optical potentials requires to go beyond order-by-order addition of terms and to aim for a stochastic resummation by sampling a limited number of representative higher order diagrams.
Figure 1: Approximate number of Feynman diagrams contributing to the self-energy for two and three-body forces. The values shown in green give the exact number of Feynman diagrams for the first three orders [28] while valuse at higher order are estimated based on their factorial scaling. For each order, very few representative diagrams are shown.
Sampling different diagram topologies is the fundamental idea of the diagrammatic Monte Carlo (DiagMC) approach  [29]. This method has first gained much interest in the field of condensed matter physics, where it was employed to study infinite systems at finite temperature such as the polaron [30] and the unitary Fermi gas [31, 32]. More recent advances in the DiagMC framework have been applied to study single-particle excitations in the electron gas [33, 34] and the real-time dynamics of dissipative quantum impurity models [35]. To our knowledge, DiagMC has still been poorly investigated for bound systems at zero temperature such as atomic nuclei, where (large) discrete bases are usually needed. Furthermore, applications to scattering theory require satisfying the causality principle, which is encoded in a dispersive spectral relation. This property is naturally broken by sharp truncations in the Feynman perturbation series, and it has never been enforced in previous DiagMC calculations. In this Letter, we introduce a DiagMC algorithm for the Richardson model and impose the correct dispersive relation by representing the stochastic self-energy with a parameterization suitable for finite systems.
The Richardson model was introduced as a simplified description of nuclear superfluidity. It explains pairing correlations in Sn and Pb isotopes with considerable accuracy [36, 37], as well as in quantum chemistry systems [38]. Furthermore, it provides stringent tests for quantum many-body theories that aim to describe strongly correlated Fermi liquids [39], and has challenged quantum neural networks [40] and eigenvector continuation [41]. The availability of exact solutions makes it suitable for precise benchmarks.
The Richardson Hamiltonian includes a contact interaction between nucleons that occupy the same energy level [42]. We consider $D$ doubly degenerate and equidistant energy levels and we focus on the half-filled system with $A=2P$ fermions,
$H^{(D)}=\xi\sum_{p=1}^{D}\sum_{s=\uparrow,\downarrow}\left(p-1\right)c_{ps}^{\dagger}c_{ps}-\frac{g}{2}\sum_{p,q=1}^{D}c_{p\uparrow}^{\dagger}c_{p\downarrow}^{\dagger}c_{q\downarrow}c_{q\uparrow},$ (1)
where we set $\xi=1$ without loss of generality. The ground state of this Hamiltonian can be written analytically as the product of the $P$ lowest correlated-pair creation operators, $b^{\dagger}_{i}$, such that |$\ket{\Psi^{A}_{0}}=\prod_{i=1}^{P}b_{i}^{\dagger}\ket{0}$ [43].
The propagator theory of the Richardson model was explored in Refs. [27] and [44]. The one-body propagator is diagonal and invariant under spin exchange, and hence only its $D$ diagonal and spin-up components need to be considered. Furthermore, we found that ladder diagrams greatly dominate the self-energy expansion of the propagator in the perturbative regime. We will focus on this subset of diagrams in our sampling and show that this is sufficient to compute the full correlation energy within the stochastic error and the error induced by a finite regulator.
a)
b)
c)
d)
Figure 2: Updates used to sample the entire ladder diagrammatic expansion [43]. Rules a), b) and d) yield the standard Monte Carlo evaluation of single diagrams. Rule c) allows to span the topological space of all self-energy ladder diagrams.
Diagrammatic Monte Carlo. Standard Monte Carlo integration has been used to compute many-body perturbation theory (PT) for nucleonic matter reaching up to the full fourth order [45] but relies on the cumbersome determination and evaluation of every single PT contribution [46]. The DiagMC framework used in this work overcomes this limitation by promoting diagram topologies on the same footing as internal quantum numbers and frequencies  [31, 29]. We express the exact self-energy according to the fundamental identity
[…]
where $\mathcal{D}_{p}$ is a diagram characterized by its topology and internal single-particle quantum numbers and frequencies, which we collectively label $\mathcal{C}$, and $Z_{p}$ is a normalization factor. The integral over $d\mathcal{C}$ includes the sum over all topologies. If the diagrams are sampled according to a probability distribution $\propto\lvert\mathcal{D}_{p}\left(\omega;\mathcal{C}\right)\rvert$, the self-energy can be calculated as a stochastic sum of the phases of the diagrams. The sampling of ladder diagrams is implemented through a Markov chain in the space of diagrams. At each step of the chain, one of the updates in Fig. 2 is called at random, and a walker in $\mathcal{C}$-space is generated using the Metropolis-Hastings algorithm [47, 48]. We refer to the Supplemental Material for a comprehensive overview of the present DiagMC implementation [43].
The DiagMC algorithm is built upon the Feynman perturbation expansion which, if truncated sharply at each order, can break important non-perturbative features of the exact self-energy. Most notably, the causality principle that is important for scattering events may require full all-order summations. Causality is expressed through the boundary condition of the Källén-Lehmann representation for the exact self-energy:
$\Sigma_{p\uparrow}^{\star}\left(\omega\right)=\Sigma_{p}^{(\infty)}+\sum_{n=1}\frac{A_{p}^{n}}{\omega-B_{p}^{n}+i\eta}+\sum_{k=1}\frac{A_{p}^{k}}{\omega-B_{p}^{k}-i\eta},$ (3)
with $A_{p}^{n/k}>0$. Note that the dispersion relation implies that only the static mean-field term, $\Sigma_{p}^{(\infty)}$, and the imaginary part of (3) are required to reconstruct the entire self-energy [25]. The algebraic diagrammatic construction (ADC) that is state-of-the-art for nuclear physics and quantum chemistry was developed with the specific aim of retaining this structure exactly [49, 27, 50].
The $\Sigma_{p}^{(\infty)}$ term is a tadpole diagram that plays a special role in the DiagMC algorithm. Its expression includes a convergence factor $e^{i\omega\eta}$ that makes the Monte Carlo integration inefficient; however, it can be easily calculated with standard numerical methods. Moreover, DiagMC calculates the self-energy only up to the overall (real) factor $Z_{p}$ of Eq. (2). To fix this ambiguity and at the same time address the convergence problem of the tadpole diagrams, we approached the problem in a novel way. First, the $\Sigma_{p}^{(\infty)}$ diagram, and any diagram that includes repetitions of it, are simple enough that they can be evaluated exactly with an analytical frequency integration followed by standard integration over the remaining single-particle quantum numbers. Hence, we compute it separately to avoid any Monte Carlo instability. Second, the Markov chain walker of DiagMC can still be implemented to retain all steps over the tadpole diagrams: simply such contributions are ignored in the sum of Eq. (2) to avoid any double counting. This fact gives us the additional freedom needed to determine the overall normalization. We replace the tadpoles diagrams with unphysical ones that have an exactly known analytical solution and compare it with the partial DiagMC resummation over the same terms. We can deduce the overall normalization factor by counting how many times the unphysical diagram is visited compared to the total number of samples [43].
Within the radius of convergence for the perturbation expansion, an increase in the truncation order must bring the simulation closer to the correct Källén-Lehmann representation. An important novelty in our approach is to perform a nonlinear least-square minimization (NLLSM) that fits the imaginary part of the self-energy calculated with DiagMC as a sum of Lorentzians. Dispersion relations are used to recover a natural parameterization of the full self-energy that fulfills the causality principle. This approach allows to control the Monte Carlo noise of the simulation by translating it into a (small) error on the fitted parameters [43].
Results.
We perform calculations in the so-called $sc0$ framework [51, 52]. This is a standard prescription in nuclear theory that consists in imposing a fully self-consistent calculation for the static component of the self-energy, $\Sigma_{p}^{(\infty)}$, while the dynamic part is expanded with respect to a mean-field reference propagator. Fig. 3 reports the components $p=1$ and $p=6$ of the imaginary part of the irreducible self-energy for $g=0.3$, $\eta=0.1$, and $D=10$. The results were obtained with a Hartree-Fock (HF) reference propagator by sampling ladder diagrams up to order eight. The simulated self-energy shows only small deviations from the exact analytic form, Eq. (3), so that the NLLSM fit is extremely precise.
Figure 3: Imaginary part of the diagonal components $p=1,6$ of the irreducible self-energy. The self-energy was calculated with DiagMC and fitted via NLLSM with a sum of Lorentzian functions. The simulation was performed with a HF reference propagator in a model space of dimension $D=10$ and $A=10$ particles. We used $g=0.3$, $\eta=0.1$. DiagMC sampled the ladder series up to the eighth order. The Lorentzian representation is very accurate due to the small stochastic noise, which is barely visible only in the far tails.
Our implementation of DiagMC carries three sources of uncertainty. The first is the statistical error that is intrinsic to any Monte Carlo approach and can be controlled to a good extent by additional sampling. Second, we do perform a many-body truncation by restricting our diagram space up to eighth order in PT and to ladder diagrams only. The error involved is negligible because excluded diagram topologies are known to be irrelevant for the Richardson Hamiltonian [27] and the first eight PT orders are substantially converged. The comparison between the NLLSM fit of Eq. (3) and the DiagMC sampling, shown in Fig. 3, demonstrates the quality of our simulations. Finally, we have a dependence on the regulator parameter $\eta$ [43]. In principle, the exact result is recovered in the limit $\eta\to 0$, where it would cause numerical instabilities. In fact, however, a finite value of $\eta$ simply sets the energy resolution of our simulations, and how ‘small’ it should be depends on the specific application.
Figure 4: Correlation energy for different values of the regulator $\eta$. The figure shows results for the coupling constants $g=0.1$ and $g=0.3$. The solid lines are the full configuration interaction (FCI) results, and they are obtained from the exact diagonalization of the Richardson Hamiltonian.
To study the dependence of our results on $\eta$ we performed simulations using different regulator values in the model space $D=10$ for $g=0.1$ and $g=0.3$. Fig. 4 shows a moderate dependence on the regulator $\eta$. The deviation between our calculations and the true correlation energies is at most 1% for both $g=0.1$ and $g=0.3$. Hence, we take $1\%$ as the error for all our calculations of correlation energies and always use $\eta=0.1$.
Figure 5: Correlation energy of the Richardson model at half-filling. The regulator is set at $\eta=0.1$ and the error of $1\%$ on the correlation energy is too small to be visible in the plot. Top panel: Dependence of the correlation energy on the coupling constant $g$ in the model space $D=10$. Bottom panel: Correlation energy per particle for different model space dimensions and $g=0.3$.
The correlation energies derived from the Migdal-Galitskii-Koltun (MGK) sum rule [53, 54] are benchmarked against the exact full configuration interaction (FCI) solutions and the ADC(3) many-body truncation values in Fig. 5. The top panel shows the correlation energy dependence on the coupling constant $g$ within the model space of size $D=10$. The bottom panel shows the correlation energy per particle with changing model space size $D$, at half-filling, and $g=0.3$. We find agreement with the exact energies within the error bars of the Monte Carlo uncertainty, confirming that the present ladder resummation dominates the self-energy expansion. Remarkably, DiagMC with the HF reference propagator outperforms ADC(3). This discrepancy is to be ascribed to the presence of time inversion terms in the ladder series, which is resummed to all orders in ADC(3) but only in a Tamm-Dancoff fashion [27]. DiagMC correctly resums the Feynman ladders, accounting for all possible time orderings.
Conclusion. For the weak-coupling regime, we conclude that DiagMC on bound systems (with discrete spectra) can match and even supersede state-of-the-art methods for computing many-body propagators, including the ADC(3) truncation scheme employed in ab initio nuclear physics and quantum chemistry. The Richardson model could be solved without the need to exploit convegence acceleration methods, such as the Borel resummation that has already been widely used by DiagMC applications to condensed matter [55, 56]. This leaves room for approaching more realistic problems. It should be stressed that this work focused on the Green’s function; however, DiagMC can be extended to other methods based on diagrammatic expressions, such as coupled cluster amplitudes or many-body PT contributions. Therefore, DiagMC holds significant potential for advancing various fields of physics involving bound many-fermion systems at zero temperature. The success of the present calculations signifies an initial step towards a unified description of nuclear structure and reactions by performing a stochastic sampling over the space of Feynman diagrams of a simplified pairing Hamiltonian.